OpenStudy (anonymous):

Let $ defined for all nonzero integers a and b such that : a$b = 2a=2b/2ab What is the vaule of 2$3? Thanks ahead of time :]

OpenStudy (anonymous):

Also another part to this question : What is the value of c if 2.4c=96

OpenStudy (anonymous):

c = 40

OpenStudy (anonymous):

Got that one. need help for the top one

OpenStudy (anonymous):

I got 10/12

OpenStudy (anonymous):

a$b = 2a = 2a/2ab so 2$3 = 2*2 = 2*2/2*2*3

OpenStudy (anonymous):

Ohh error on top so sorry 2a+2b/2ab

OpenStudy (anonymous):

I forgot to push shift =___=

OpenStudy (anonymous):

\[a $ b = 2a+\frac{2b}{2ab}\] \[a$b = 2a+\frac{1}{a}\]

OpenStudy (anonymous):

there is no "b" on the right hand side. so \[2$x = 2\times 2+\frac{1}{2}=4.5\]